OpenStudy (anonymous):

Four times the difference between 44 and Sarah's age (a) is 64. Assume Sarah is younger than 44. Represent the above statement by an equation.

pooja195 (pooja195):

Start out by writing the problem mathematically. Let Sarah's age = x Therefore, four times the difference between 44 and x is 64 can be written 4x-44 = 64 Now solve for x. Add 44 to both sides. 4x = 108 Divide each side by 4. x = 27. Therefore Sarah is 27 years old.

pooja195 (pooja195):

i added the solve for x just in case if it asked

pooja195 (pooja195):

Welcome to Openstudy! :)

OpenStudy (anonymous):

is this the answer 4(a-44)=64

pooja195 (pooja195):

yes i beleive thats right

OpenStudy (anonymous):

A salesman gets $40 per day as wages and $4.50 as commission for every pair of athletic shoes he sells in a day. Write an equation to represent how many pairs of shoes p he must sell in a day to make $112 on that day.

OpenStudy (anonymous):

can u help me with this one 2

pooja195 (pooja195):

sure :)

pooja195 (pooja195):

: 4.50P + 40 = 112

OpenStudy (radar):

There was an error for the first problem, 27 is not the correct answer.....but real close.

OpenStudy (radar):

thatbeautygal101 expressed the equation correctly. 4( a - 44) = 64

OpenStudy (anonymous):

John has read the first 114 pages of a novel. He has read 3 pages less than one-third of the novel. Write an equation to describe the total number of pages p in the novel, and find the value of p. hellppp

OpenStudy (radar):

114 + 3 = p/3 117 = p/3 both sides times 3 getting p = 351

OpenStudy (radar):

Sara's age: 4(44 - a) = 64 Divide both sides by 4 getting 44-a = 16. Subtract 44 from both sides getting: -a = 16 - 44 or -a = -28 Multiply both sides by -1 getting a = 28.